OpenStudy (anonymous):

3. Write the point-slope form of the equation passing through (5, -1) with a slope of 6.(Points : 1) y-1=6(x+5) y+5=6(x-1) y+1=6(x-5) y-5=6(x+1) 4. Write the equation for the vertical line passing through (5, 8).(Points : 1) x = 5 x = 8 x = 0 x = 5/8

OpenStudy (blurbendy):

y + 1 = 6(x -5) 4) x = 5